# Geodesic loops and orthogonal geodesic chords without self-intersections

### v1

@@ 1. Results
The author has shown a perturbation result for geodesic segments in [18, Lem.3]. As a consequence we obtain for a point $p$ in a compact Riemannian manifold of dimension $n\geq 3$ with injectivity radius $\mathrm{inj}=4\eta>0$ the following statement: If $N\geq 2$ geometrically distinct geodesic segments $c_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,N$ parametrized by arc length intersect in the common midpoint $p=c_{j}(0)$ then one can perturb away the intersection of the geodesic segments. For a precise statement see Lemma 2.
[…]
@@
Instead of considering closed resp. periodic geodesics one can study different boundary conditions for geodesic segments. In this paper we study first for points $p,q\in M$ geodesic segments $c:\mathcal{I}\longrightarrow M$ joining $p=c(0)$ and $q=c(1).$ Here we use as parameter space the unit interval $\mathcal{I}=[0,1]$ if $p\not=q$ and the circle $\mathcal{I}=S^{1}=\mathbb{R}/\mathbb{Z}=[0,1]/\{0,1\}$ in case of a geodesic loop, i.e. $p=q.$ We call the geodesic segment $c:\mathcal{I}\longrightarrow M$ simple, resp. without self-intersection, if the map $c:\mathcal{I}\longrightarrow M$ is injective. We call two geodesic segments $c_{1},c_{2}:\mathcal{I}\longrightarrow M$ geometrically distinct, if $c_{1}(\mathcal{I})\not=c_{2}(\mathcal{I}).$ We say that two geometrically distinct geodesic segments $c_{1},c_{2}:\mathcal{I}\longrightarrow M$ do not have an interior intersection, if $c_{1}((0,1))\cap c_{2}((0,1))=\emptyset.$ It follows from a general result by Bettiol and Giambò presented in  [2, Theorem 5.10] that for a generic Riemannian metric on a compact manifold $M$ the geodesic segments, resp. geodesic loops, are non-degenerate in the sense of Morse theory. We show in this paper that we can use the perturbation result Lemma 2 and the genericity result of [2, Theorem 5.10] to show the following result:
@@
Note that this also shows that for points $p,q$ for a generic metric there is no closed geodesic passing through $p$ and $q.$ It even shows that for a generic metric any geodesic through $p$ and $q$ does not go through $p$ resp. $q$ again. In the particular case $p=q$ it follows from Theorem 1 that all geodesic loops for a generic metric are prime, i.e. there is no shorter geodesic loop, which is a part of the given geodesic loop, cf. [16]. Hence for a generic metric for any two geodesics $c,d:[0,1]\longrightarrow M,c(0)=d(0)=p,c(1)=d(1)=q$ we have either $c((0,1))=d((0,1))$ or $c([0,1])\cap d([0,1])=\{p,q\}.$
@@
Let $M$ be a simply-connected and compact manifold of dimension $n\geq 3$ and $p,q\in M.$ For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all geodesic segments joining $p$ and $q$ are non-degenerate and simple, and geometrically distinct ones do not have interior intersection. For the number $N_{g}(p,q;t)$ of geodesic segments joining $p$ and $q$ of length $\leq t$ we obtain $\liminf_{t\to\infty}(N_{g}(p,q;t)/t)>0.$
This gives a quantitative version of Serre’s result [20, Ch.IV.,Sec.7, Prop.14] which also gives a geometrically satisfying result concerning geometrically distinct geodesic segments between two fixed points, cf. the discussion in the doctoral thesis by Mentges [16, Sec.1.2] or by Tanaka [21, Problem B]. Another boundary condition for geodesic segments is the following: Let $(M,g)$ be a compact Riemannian manifold with smooth boundary $\partial M$ and interior $\textrm{Int}(M)=M-\partial M$ then a geodesic chord $c:[0,1]\longrightarrow M$ is a geodesic segment with $c(0),c(1)\in\partial M$ and $c(t)\in\textrm{Int}(M)$ for all $t\in(0,1).$ It is called an orthogonal geodesic chord if $c$ meets the boundary $\partial M$ orthogonally. It is called an orthogonal-tangent geodesic chord if it is tangent to the boundary at $c(0)$ and orthogonal to the boundary at $c(1).$ As a consequence of the genericity result by Bettiol and Giambò [2] we show:
@@
Lusternik and Schnirelmann [14] have shown that a bounded domain in $\mathbb{R}^{n},n\geq 2$ with a smooth and convex boundary has at least $n$ double normals, i.e. lines intersecting the boundary orthogonally. This was generalized for a Riemannian metric, Bos [4, p.431] showed that on an $n$-disc $D^{n}$ with a Riemannian metric whose boundary $\partial D^{n}$ is strictly convex there are $n$ geometrically distinct orthogonal geodesic chords. Using a generalization of this result due to Giambò, Giannoni, and Piccione [7, Theorem 1.1] we obtain from Theorem 2:
@@
For $n\geq 3$ let $\mathcal{G}(D^{n})$ be the space of Riemannian metrics on the $n$-disc $D^{n}$ with the strong $C^{k}$-topology with $k\geq 2.$ The subset $\mathcal{G}_{1}=\mathcal{G}_{1}(D^{n})$ of Riemannian metrics on $D^{n}$ for which there is no orthogonal-tangent geodesic chord, is an open and non-empty subset of $\mathcal{G}(D^{n}).$ Then there is an open subset $\mathcal{G}_{2}\subset\mathcal{G}_{1}$ whose closure $\overline{\mathcal{G}}_{2}$ coincides with the closure $\overline{\mathcal{G}}_{1},$ i.e. $\overline{\mathcal{G}}_{1}=\overline{\mathcal{G}}_{2},$ such that for all $g\in\mathcal{G}_{2}$ there are $n$ geometrically distinct geodesic orthogonal chords which are non-degenerate and have no self-intersection and do not intersect each other.
Here we note that geometrically distinct geodesic chords cannot intersect at a boundary point. The boundary $\partial D^{n}$ is called strictly convex, if all principal curvatures $\kappa_{i},i=1,\ldots,n-1$ of the boundary with respect to the inward pointing unit normal, are positive, i.e. $\kappa_{i}>0,i=1,\ldots,n-1.$ In particular there is no geodesic in the disc touching the boundary $\partial D^{n}$ and hence there is no orthogonal-tangent geodesic chord. Therefore we obtain from Theorem 3 the following result in the spirit of Bos’ theorem [4, p.431]:
@@
There is an open and dense subset $\mathcal{G}_{*}$ of the set $\mathcal{G}_{+}(D^{n})$ of Riemannian metrics on the $n$-dics $D^{n}$ of dimension $n\geq 3$ for which the boundary $\partial D^{n}$ is strictly convex such that for all Riemannian metrics $g\in\mathcal{G}_{*}$ there are at least $n$ geometrically distinct orthogonal geodesic chords without self-intersections. And these geodesic chords do not intersect each other.
Using a curve shortening construction called disc flow on surfaces Hass and Scott [11, Thm.4.2] have shown that a Riemannian $2$-disc with strongly convex boundary carries at least two geometrically distinct orthogonal geodesic chords without self-intersections. For the proof one can also use the curve shortening flow, as shown by Ko [13]. With a curve shortening flow Li [15] presents a proof of the above cited result by Lusternik and Schnirelmann about the existence of $n$ double normals of a strictly convex compact domain in $\mathbb{R}^{n}.$
@@ 2. Intersection of geodesic segments and perturbation
For the intersection points of geodesic segments we obtain the following result which is similar to the corresponding result for closed geodesics, cf. [18, Lem.2] resp. [2, Sec.2.3]. Recall from the Introduction that we denote a geodesic segment $c:\mathcal{I}\longrightarrow M$ joining $p=c(0),q=c(1)$ using the parameter space $\mathcal{I}=[0,1]$ if $p\not=q$ and $\mathcal{I}=S^{1}$ for $p=q.$
@@
If $\mathrm{inj}>0$ is the injectivity radius of $(M,g),$ we conclude from $\gamma(t)=\gamma(s),t<s$ that $(s-t)L(\gamma)\geq\mathrm{inj}.$ This implies that for all $t\in[0,1]:\#\gamma^{-1}(\gamma(t))\leq L(\gamma)/\mathrm{inj}+1.$
(a) If the set $DP_{1}(\gamma)$ is not finite, then there exist sequences $s_{j}<t_{j}\in[0,1]$ converging to $s^{*},t^{*}\in[0,1]$ with $s_{j}\not=s^{*},t_{j}\not=t^{*},(t_{j}-s_{j})L(\gamma)\geq\mathrm{inj}$ for all $j,$ $(t^{*}-s^{*})L(\gamma)\geq\mathrm{inj},$ and $\gamma(s_{j})=\gamma(t_{j})$ for all $j\geq 1.$ For $p_{1}=\gamma(t^{*})=\gamma(s^{*})$ set $v=c^{\prime}(t^{*}),w=c^{\prime}(s^{*}).$ The geodesic segment $\gamma$ is parametrized proportional to arc length, hence $\|v\|=\|w\|.$ Since $c(t_{j})=\exp_{p}((t_{j}-t^{*})v)=c(s_{j})=\exp_{p}((s_{j}-s^{*})w)$ we obtain for sufficiently large $j$ with $|t_{j}-t^{*}|\|v\|,|s_{j}-s^{*}|\|w\|<\mathrm{inj}$ that $(t_{j}-t^{*})v=(s_{j}-s^{*})w$ holds, i.e. $v=\pm w.$ If $v=w$ then $\gamma(t)=\gamma(t+(t^{*}-s^{*}))$ for all $t,$ i.e. $\gamma$ is a closed geodesic, which is not prime. If $v=-w$ we conclude $\gamma((t^{*}-s^{*})/2-t)=\gamma((t^{*}-s^{*})/2+t)$ for all $t,$ showing that $\gamma$ is not smooth at $(t^{*}-s^{*})/2.$
@@
We obtain from the perturbation result [18, Lem.3]:
@@
Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 3$ and $\eta>0$ be a positive number such that the injectivity radius $\mathrm{inj}$ satisfies $\mathrm{inj}\geq 4\eta.$ For a point $p\in M$ let $c_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,k$ be $k\geq 2$ geometrically distinct geodesic segments parametrized by arc length with common midpoint $p=c_{j}(0),j=1,\ldots,k.$
Then there is in any neighborhood of the Riemannian metric $g$ with respect to the strong $C^{r}$-topology with $r\geq 2$ a Riemannian metric $\overline{g}$ coinciding with $g$ outside the geodesic ball $B_{p}(\eta)$ of radius $\eta$ centered at $p$ carrying geodesic segments $\overline{c}_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,k$ parametrized by arc length which coincide with $c_{j}$ outside $B_{p}(\eta).$ These geodesic segments are simple and do not intersect each other.
A similar proof as in Lemma [18, Lem.3] shows the following
@@
Let $c:[0,1]\longrightarrow M$ be a geodesic loop with $p=c(0)=c(1)$ of the Riemannian metric $g$ on the differentiable manifold $M$ with $v=c^{\prime}(0)\in T_{p}M,w=c^{\prime}(1)\in T_{p}M.$ Then there exists a continuous one parameter perturbation $g_{s},s\in[0,s_{1}]\in\mathcal{G}(M)$ in the space of Riemannian metrics with the strong $C^{k}$-topology with $k\geq 2$ of the metric $g=g_{0}$ and continuous functions $s\in[0,s_{1}]\mapsto v(s),w(s)\in T_{p}M$ such that $c_{s}:[0,1]\longrightarrow M$ is a geodesic loop with $p=c(0)$ of the Riemannian metric $g_{s}$ with $v(s)=c_{s}^{\prime}(0),w(s)=c_{s}^{\prime}(1)$ and $v(s),v$ resp. $w(s),w$ are linearly independent for $s>0.$ The metric $g_{s}$ agrees with $g$ outside an arbitrarily small geodesic ball around $p.$

### v2

@@ 1. Results
It is a well known result that on any compact smooth manifold $M$ with tangent bundle $TM$ endowed with a Riemannian metric $g$ there exists a non-trivial closed geodesic. If the manifold is not simply-connected a shortest non-contractible curve is a closed geodesic. In the simply-connected case this result is due to Lusternik and Fet [17], cf. [15, Sec.2.4]. We call a closed geodesic $c:S^{1}=\mathbb{R}/\mathbb{Z}\longrightarrow M$ prime if it is not the covering of a shorter closed curve, then the velocity field $c^{\prime}:S^{1}\longrightarrow TM$ defines an injective map. We call a closed geodesic simple resp. without self-intersections if the map $c:S^{1}\longrightarrow M$ is injective. In particular a simple closed geodesic is prime. The shortest noncontractible smooth curve on a manifold is a simple closed geodesic. On one hand any closed surfaces carries a simple closed geodesic, cf. [23, Rem. 5]. On the other hand it is an open question whether any Riemannian manifold $M$ of dimension $n=\dim(M)\geq 3$ carries a simple closed geodesic, cf. [2, Introduction]. The situation is quite different for a generic metric in dimensions $\geq 3.$ The author has shown recently that that for a generic Riemannian metric, resp. a generic reversible Finsler metric, on a compact manifold of dimension $n\geq 3$ all prime closed geodesics have no self-intersections, and they do not intersect each other, cf. [23, Thm.1].
@@
One ingredient of the proof is the following perturbation result for intersecting geodesic segments: If $N\geq 2$ geometrically distinct geodesic segments $c_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,N$ parametrized by arc length intersect in the common midpoint $p=c_{j}(0)$ for sufficiently small $\eta>0$ then one can perturb away the intersection of the geodesic segments. For a precise statement see Lemma 2.
The other ingredient is the bumpy metric theorem for closed geodesics stating that for a generic Riemannian metric on a compact manifold all closed geodesics are non-degenerate in the sense of Morse theory. A closed geodesic $c:S^{1}\longrightarrow M$ is called non-degenerate if there is no non-trivial periodic Jacobi field along $c$ which is orthogonal to the velocity field $c^{\prime}.$ The statement is due to Abraham [1], proofs were given by Anosov [3], and White [27, Thm 2.2], see also the book by Moore [21, §2.7]. Hence the energy functional $E:\Lambda M\longrightarrow\mathbb{R},E(\sigma)=\frac{1}{2}\int_{0}^{1}g(\sigma^{\prime}(t),\sigma^{\prime}(t))\,dt$ on the Hilbert manifold $\Lambda M$ of closed curves $\sigma:S^{1}=[0,1]/\{0,1\}\longrightarrow M$ (called the free loop space) is a Morse-Bott function whose critical set is the disjoint union of orbits of closed geodesics under the canonical action of the orthogonal group $\mathbb{O}(2)$ in dimension $2$ acting on $S^{1}=\mathbb{R}/\mathbb{Z}=[0,1]/\{0,1\}.$ It also follows that for any $L>0$ there are only finitely many geometrically distinct closed geodesics with length $\leq L.$ It is a consequence of the perturbation result and the bumpy metric theorem that for a generic Riemannian metric on a compact manifold of dimension $n\geq 3$ all prime closed geodesics have no self-intersections, i.e. are simple, and they do not intersect each other, as shown by the author, cf. [23, Thm.1]. Note that on a surface with a non-degenerate closed geodesic with self-intersection it is not possible to perturb away the self-intersection.
@@
Instead of considering closed resp. periodic geodesics one can study different boundary conditions for geodesic segments. In this paper we study first for points $p,q\in M$ geodesic segments $c:\mathcal{I}\longrightarrow M$ joining $p=c(0)$ and $q=c(1).$ Here we use as parameter space the unit interval $\mathcal{I}=[0,1]$ if $p\not=q$ and the circle $\mathcal{I}=S^{1}=\mathbb{R}/\mathbb{Z}=[0,1]/\{0,1\}$ in case of a geodesic loop, i.e. $p=q.$ Note that in the case of a geodesic loop $c:S^{1}\longrightarrow M$ only the map $c:[0,1]\longrightarrow M$ is smooth, since in general for a geodesic loop in contrast to a closed resp. periodic geodesic we have $c^{\prime}(0)\not=c^{\prime}(1).$ We call the geodesic segment $c:\mathcal{I}\longrightarrow M$ simple, resp. without self-intersection, if the map $c:\mathcal{I}\longrightarrow M$ is injective. We call two geodesic segments $c_{1},c_{2}:\mathcal{I}\longrightarrow M$ geometrically distinct, if $c_{1}(\mathcal{I})\not=c_{2}(\mathcal{I}).$ We say that two geometrically distinct geodesic segments $c_{1},c_{2}:\mathcal{I}\longrightarrow M$ do not have an interior intersection, if $c_{1}((0,1))\cap c_{2}((0,1))=\emptyset.$ It follows from a general result by Bettiol and Giambò presented in  [4, Theorem 5.10] that for a generic Riemannian metric on a compact manifold $M$ the geodesic segments, resp. geodesic loops, are non-degenerate in the sense of Morse theory, cf. Section 3. We show in this paper that we can use the perturbation result Lemma 2 and the genericity result of [4, Theorem 5.10] to show the following result:
@@
Note that Theorem 1 also shows that for points $p,q$ for a generic metric there is no closed geodesic passing through $p$ and $q,$ cf. Lemma 6. It even shows that for a generic metric any geodesic through $p$ and $q$ does not go through $p$ resp. $q$ again. In the particular case $p=q$ it follows from Theorem 1 that all geodesic loops for a generic metric are prime, i.e. there is no shorter geodesic loop, which is a part of the given geodesic loop, cf. [20]. Hence for a generic metric for any two geodesics $c,d:[0,1]\longrightarrow M,c(0)=d(0)=p,c(1)=d(1)=q$ we have either $c((0,1))=d((0,1))$ or $c([0,1])\cap d([0,1])=\{p,q\}.$
@@
Let $M$ be a simply-connected and compact manifold of dimension $n\geq 3$ and $p,q\in M.$ For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all geodesic segments joining $p$ and $q$ are non-degenerate and simple, and geometrically distinct ones do not have interior intersection. For the number $N_{g}(p,q;t)$ of geometrically distinct and simple geodesic segments joining $p$ and $q$ of length $\leq t$ we obtain $\liminf_{t\to\infty}(N_{g}(p,q;t)/t)>0.$
This is a quantitative version of Serre’s result [25, Ch.IV., Sec.7, Prop.14] giving a result concerning geometrically distinct geodesic segments between two fixed points. Therefore Corollary 1 is also satisfying from a geometric point of view, cf. the discussion in the doctoral thesis by Mentges [20, Sec.1.2] or by Tanaka [26, Problem B]. On a surface, i.e. for $n=2,$ we cannot expect to remove intersections of geodesic segments by a small perturbation. But for a point $p$ and a generic metric on a closed surface, all geodesic loops based at $p$ are prime, cf. Proposition 1.
Another boundary condition for geodesic segments is the following: Let $(M,g)$ be a compact Riemannian manifold with smooth boundary $\partial M$ and interior $\textrm{Int}(M)=M-\partial M$ then a geodesic chord $c:[0,1]\longrightarrow M$ is a geodesic segment with $c(0),c(1)\in\partial M$ and $c(t)\in\textrm{Int}(M)$ for all $t\in(0,1).$ It is called an orthogonal geodesic chord if $c$ meets the boundary $\partial M$ orthogonally. Orthogonal geodesic chords are also called free boundary geodesics, cf. for example [16] or [28]. A chord $c:[0,1]\longrightarrow M$ is called an orthogonal-tangent geodesic chord if it is tangent to the boundary at $c(0)$ and orthogonal to the boundary at $c(1).$ As a consequence of the genericity result by Bettiol and Giambò [4] we show:
@@
Lusternik and Schnirelmann [18] have shown that a bounded domain in $\mathbb{R}^{n},n\geq 2$ with a smooth and convex boundary has at least $n$ double normals, i.e. lines intersecting the boundary orthogonally. This was generalized for a Riemannian metric, Bos [6, p.431] showed that on an $n$-disc $D^{n}$ with a Riemannian metric whose boundary $\partial D^{n}$ is strictly convex there are $n$ geometrically distinct orthogonal geodesic chords. Using a generalization of this result due to Giambò, Giannoni, and Piccione [9, Theorem 1.1] we obtain from Theorem 2:
@@
For $n\geq 3$ let $\mathcal{G}(D^{n})$ be the space of Riemannian metrics on the $n$-disc $D^{n}$ with the strong $C^{k}$-topology with $k\geq 2.$ Let $\mathcal{G}_{1}=\mathcal{G}_{1}(D^{n})$ be the set of Riemannian metrics on $D^{n}$ for which there is no orthogonal-tangent geodesic chord. Then $\mathcal{G}_{1}$ is an open and non-empty subset of $\mathcal{G}(D^{n}).$
Furthermore there is an open subset $\mathcal{G}_{2}\subset\mathcal{G}_{1}$ whose closure $\overline{\mathcal{G}}_{2}$ coincides with the closure $\overline{\mathcal{G}}_{1},$ i.e. $\overline{\mathcal{G}}_{1}=\overline{\mathcal{G}}_{2},$ with the following property: For all Riemannian metrics $g\in\mathcal{G}_{2}$ there are $n$ geometrically distinct geodesic orthogonal chords, which are non-degenerate and have no self-intersection, and do not intersect each other.
Here we note that geometrically distinct geodesic chords cannot intersect at a boundary point. The boundary $\partial D^{n}$ is called strictly convex, if all principal curvatures $\kappa_{i},i=1,\ldots,n-1$ of the boundary with respect to the inward pointing unit normal, are positive, i.e. $\kappa_{i}>0,i=1,\ldots,n-1.$ In particular there is no geodesic in the disc touching the boundary $\partial D^{n}$ and hence there is no orthogonal-tangent geodesic chord. Therefore we obtain from Theorem 3 the following result in the spirit of Bos’ theorem [6, p.431]:
@@
There is an open and dense subset $\mathcal{G}_{*}$ of the set $\mathcal{G}_{+}(D^{n})$ of Riemannian metrics on the $n$-dics $D^{n}$ of dimension $n\geq 3,$ for which the boundary $\partial D^{n}$ is strictly convex, with the following property: For all Riemannian metrics $g\in\mathcal{G}_{*},$ there are at least $n$ geometrically distinct orthogonal geodesic chords without self-intersections.
Furthermore, these geodesic chords do not intersect each other.
Using a curve shortening construction called disc flow on surfaces, Hass and Scott [13, Thm.4.2] have shown that a Riemannian $2$-disc with strongly convex boundary carries at least two geometrically distinct orthogonal geodesic chords without self-intersections. For the proof one can also use the curve shortening flow, as shown by Ko [16]. With a curve shortening flow, Li [19] presents a proof of the above cited result by Lusternik and Schnirelmann about the existence of $n$ double normals of a strictly convex compact domain in $\mathbb{R}^{n}.$
Acknowledgement. I am grateful to the referee for the helpful comments and suggestions.
@@ 2. Intersection of geodesic segments and perturbation
For the intersection points of geodesic segments we obtain the following result which is similar to the corresponding result for closed geodesics, cf. [23, Lem.2] resp. [4, Sec.2.3]. Recall from the Introduction that we denote a geodesic segment $c:\mathcal{I}\longrightarrow M$ joining $p=c(0),q=c(1)$ using the parameter space $\mathcal{I}=[0,1]$ if $p\not=q$ and $\mathcal{I}=S^{1}=[0,1]/\{0,1\}$ for $p=q.$
@@
Let $\mathrm{inj}>0$ be the injectivity radius of $(M,g).$ We conclude from $\gamma(t)=\gamma(s)$ and $t<s$ that $(s-t)L(\gamma)\geq\mathrm{inj}.$ This implies that for all $t\in[0,1]:\#\gamma^{-1}(\gamma(t))\leq L(\gamma)/\mathrm{inj}+1.$
(a) If the set $DP_{1}(\gamma)$ is not finite, then there exist sequences $s_{j}<t_{j}\in[0,1]$ converging to $s^{*},t^{*}\in[0,1]$ with $s_{j}\not=s^{*},t_{j}\not=t^{*},(t_{j}-s_{j})L(\gamma)\geq\mathrm{inj}$ for all $j,$ $(t^{*}-s^{*})L(\gamma)\geq\mathrm{inj},$ and $\gamma(s_{j})=\gamma(t_{j})$ for all $j\geq 1.$ For $p_{1}=\gamma(t^{*})=\gamma(s^{*})$ set $v=c^{\prime}(t^{*}),w=c^{\prime}(s^{*}).$ The geodesic segment $\gamma$ is parametrized proportional to arc length, hence $\|v\|=\|w\|.$ Since $c(t_{j})=\exp_{p}((t_{j}-t^{*})v)=c(s_{j})=\exp_{p}((s_{j}-s^{*})w)$ we obtain for sufficiently large $j$ with $|t_{j}-t^{*}|\|v\|,|s_{j}-s^{*}|\|w\|<\mathrm{inj}$ that $(t_{j}-t^{*})v=(s_{j}-s^{*})w$ holds, i.e. $v=\pm w.$ If $v=w$ then $\gamma(t)=\gamma(t+(t^{*}-s^{*}))$ for all $t,$ i.e. $\gamma$ is a closed geodesic, which is not prime. If v=wv=-w we conclude $\gamma((t^{*}-s^{*})/2-t)=\gamma((t^{*}-s^{*})/2+t)$ for all $t,$ showing that $\gamma$ is not smooth at $(t^{*}-s^{*})/2.$
@@
We obtain from the perturbation result [23, Lem.3]:
@@
Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 3$ and $\eta>0$ be a positive number, such that the injectivity radius $\mathrm{inj}$ satisfies $\mathrm{inj}\geq 4\eta.$ For a point $p\in M$ let $c_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,k$ be $k\geq 2$ geometrically distinct geodesic segments parametrized by arc length with common midpoint $p=c_{j}(0),j=1,\ldots,k.$
In any neighborhood $\mathcal{U}\ni g$ with respect to the strong $C^{k}$-topology with $k\geq 2,$ there is a Riemannian metric $\overline{g}\in\mathcal{U}$ coinciding with $g$ outside the geodesic ball $B_{p}(\eta)$ of radius $\eta$ centered at $p,$ carrying geodesic segments $\overline{c}_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,k$ parametrized by arc length, which coincide with $c_{j}$ outside $B_{p}(\eta).$
These geodesic segments are simple and do not intersect each other.
A similar proof as in Lemma [23, Lem.3] shows the following
@@
Let $c:[0,1]\longrightarrow M$ be a geodesic loop with $p=c(0)=c(1)$ of the Riemannian metric $g$ on the differentiable manifold $M$ with $v=c^{\prime}(0)\in T_{p}M,w=c^{\prime}(1)\in T_{p}M.$ Then there exists a continuous one parameter perturbation $g_{s},s\in[0,s_{1}]\in\mathcal{G}(M)$ in the space of Riemannian metrics with the strong $C^{k}$-topology with $k\geq 2$ of the metric $g=g_{0},$ and smooth functions $s\in[0,s_{1}]\mapsto v(s),w(s)\in T_{p}M,$ such that $c_{s}:[0,1]\longrightarrow M$ is a geodesic loop with $p=c(0)$ of the Riemannian metric $g_{s}$ with $v(s)=c_{s}^{\prime}(0),w(s)=c_{s}^{\prime}(1),$ and $v(s),v$ (resp. $w(s),w$) are linearly independent for $s>0.$ The metric $g_{s}$ agrees with $g$ outside an arbitrarily small geodesic ball around $p.$